8. Inspection costs at one of Iuliano Corporation's factories are listed below: Units produced Inspection cost...

8.

Inspection costs at one of Iuliano Corporation's factories are listed below:

Units produced Inspection cost
  February 915     $ 17,212   
  March 969     $ 17,600   
  April 921     $ 17,265   
  May 905     $ 17,110   
  June 927     $ 17,294   
  July 912     $ 17,180   
  August 929     $ 17,332   
  September 869     $ 16,660   
  October 908     $ 17,138   
Management believes that inspection cost is a mixed cost that depends on units produced.

Using the high-low method, the estimate of the fixed component of inspection cost per unit produced is closest to: (Round the intermediate calculations to two decimal places.)
Answer #1

Solution

Variable cost per employee $                      9.40
Fixed cost $              8,491.40

Working

Cost No. of activities
A High Level $          17,600.00 969
B Low Level $          16,660.00 869
C=A-B Difference $                940.00 $                  100.00
A Cost difference $                940.00
B No. of activities difference 100
C=A/B Variable cost per unit $                     9.40
High Level Low Level
A No. of activities 969 869
B Variable cost per unit $                9.4000 $                  9.4000
C=AxB Total Variable cost $             9,108.60 $              8,168.60
D Total cost $          17,600.00 $            16,660.00
E=D-C Total fixed cost $             8,491.40 $              8,491.40